# Black Big Boxes: Do Language Models Hide a Theory of Adjective Order?

### v1

@@ 7 Experiment 4: The Role of Context
We compute the expectation over a random sample of 500 contexts, taken from CAP ($\mathcal{C}$). To express the relative impact of adding context, we regress out a model’s isolated AOP-$\Delta$ score by subtracting it from the contextual scores, which we call $c\textup{-}\Delta$:
By computing this score for both the original context and for the expectation over random contexts we can quantify how much more important the original context is for AOP-$\Delta$.
We plot the results for this experiment in Figure 6B. The true context improves isolated AOP-$\Delta$ in 65.1% of the cases, whereas random contexts lead to an improvement for only 39.6%. Furthermore, for 33.2% the true context leads to an improvement whereas random contexts don’t: this subset can serve as a valuable starting point for investigating what properties of the context are important for adjective order.
Conclusions
AOP in Pythia is context-sensitive. Providing the model with the sentence where the adjective combination appears increases the model’s preference towards the target adjective order – both compared to the noun phrase in isolation and to the noun phrase in a random context. This result points in the direction of a more linguistically interesting mechanism for AOP than reliance on $n$-gram counts in training data alone. In particular, it shows a non-trivial interaction between the noun phrase and bigger context that turns out relevant for adjective placement. Establishing this opens up a possibility of studying the forms and types of this contextual interaction – a potential novel contribution to linguistic theories of adjective order.
We leave a systematic discovery of types of AOP context-dependence open for future work, and focus on a small-scale qualitative study for now. An inspection of context-sensitive cases from CAP reveals at least two ways in which context matters for AOP:
1. Structural parallelism
The ‘canonical’ word order can be overridden when context contains a parallel noun phrase with adjectives in non-canonical order (the example shows contextual AOP according to Pythia-1.4b):
  • a)
    The green big cars AOP-$\Delta$ 4.2
  • b)
    The story is not about small red houses, it’s about big green cars AOP-$\Delta$ 6.63
2. Property saliency
The property salient in the context brings the adjective describing this property to the left edge of the noun phrase. If the sentence is about color, color adjective will come first; if the sentence is about size, it will bring size adjectives to the left (again, orders preferred by Pythia-1.4b):
  • a)
    They asked my favourite color of houses and I said I liked red small houses AOP-$\Delta$ 0.32
  • b)
    They asked my favourite size of houses and I said I liked small red houses AOP-$\Delta$ 1.61
These observations, of course, barely scratch the surface of possible contextual AOP interactions, but exemplify the power of LMs in incorporating contextual information into word order judgments.
@@ 8 Discussion
Abstract Notions of Adjective Order
In this paper, we looked for AOP generalization in LMs. We were interested in whether abstract driving forces are at play in LM predictions beyond simple $n$-gram counts in training data. From our experiments we conclude that LMs show limited AOP generalization. Here, we want to highlight an important question that is rarely brought up in linguistic generalization discussions in NLP: How general and abstract is human linguistic behavior?
Linguistic literature suggests that human language production, including word order preferences, is guided by a mix of memorization and abstract principles (O’Donnell, 2015). Speakers tend to recycle more routinized and therefore more accessible combinations (MacDonald, 2013), guided by sequence frequencies in linguistic input that, in turn, form sequential associations that lead to production automatizing (Bybee, 2010; Diessel, 2019). This is compatible with a picture of language acquisition where knowledge of grammar is formed based on encountered exemplars of usage, which are stored in memory (Bybee, 2010) and serve as basis of generalization (Goldberg, 2006), but can also serve as patterns for routinized language production. So, human linguistic knowledge and language production is not totally abstract and is not based exclusively on generalization. This has been discussed across linguistic phenomena and has been suggested for word order (MacDonald, 2013; Levshina et al., 2023), although not for adjective order specifically.
When investigating generalization in LMs it is therefore important to make a distinction between the degree to which they create generalizations for a particular phenomenon, and how similar this generalization is to human behavior. In principle, one could even find cases of LM overgeneralization, if we are interested in how human-like LMs are linguistically. Importantly, in order to answer the question of human-like generalization, we would need data on the trade-off in memorization and generalization in human linguistic behavior. Such data does not exist yet for adjective order preferences.
@@
Despite the impressive performance of LMs in recent years, their impact on theoretical linguistics has remained minimal (Baroni, 2022). There are various reasons for this. On the one hand, the cognitive plausibility of current LMs remains questionable, both in terms of the scale at which they are trained (Huebner et al., 2021; Warstadt et al., 2023) and their inductive biases that depend strongly on model architecture and are not reflective of human language processing (Wilson and Frank, 2023; Oh et al., 2024). On the other hand, their black box nature makes it highly challenging to extract linguistic patterns that can be tested against existing linguistic theories. While advancements in interpretability have made this latter point less problematic, the number of works explicitly connecting LM behavior to human language processing remains few and limited to simple syntactic phenomena (e.g., Lakretz et al., 2021).
In this paper, we explore the possibility that LMs may ‘hide’ a theory of adjective order, by encoding features that are not covered by linguistic theories on adjective order. We show that LMs far exceed the predictive power of prior cognitive predictors, which raises the question whether this increase stems from meaningful interpretable features. While a lot of the LM predictions can be explained by collocational patterns—and those statistics alone do not explain the distribution of adjective orders—we do find some generalization to unseen adjective combinations and a linguistically meaningful pattern in the way that context boosts AOP. We hope our work can serve as an inspiration for using LMs as tools for measuring context-dependence, which is hard to quantify with traditional quantitative methods.
Corpus Interventions
Our analysis of AOP abstractness is constrained by the fixed training corpus the models are trained on. More detailed insights into this question can be obtained by a intervention-based approach, in which we filter out specific constructions from the training data (Jumelet et al., 2021; Misra and Mahowald, 2024; Patil et al., 2024). An interesting experiment in this direction would be to remove all sentences containing double adjectives from the training corpus, and measure whether the remaining distribution of adjectives contains a sufficient signal to acquire adjective order constraints.
Limitations
Our analysis only focuses on a single language model family (Pythia), in a a single language (English). To test the cross-lingual properties of adjective order, a wider range of language models should be examined. Our double adjective extraction method (§3.2) depends on the quality of the dependency parser that was used, and as such contains inaccuracies. Furthermore, a downside of using such a broad-coverage approach is that it becomes harder to draw targeted conclusions about the linguistic properties of cases on which the model fails or succeeds. Annotating cases containing properties that have been shown to be important for adjective order in the linguistics literature would provide deeper insights into model behavior on a linguistic level.
@@ Appendix A Complementary Results
[…]
We provide a sample of CAP examples in Table 2, categorised by the contextual and isolated AOP scores of Pythia-12b.
[…]
CAP order always preferred
1. In solution, lead(II) hydroxide is a somewhat weak base, forming lead(II) ion, Pb, under weakly acidic conditions. $16.5$ $18.8$
2. It has the only deep water port in Cambodia. $16.9$ $15.4$
3. The greatest social reforms in Denmark are certainly the work of the last half-century. $15.4$ $16.3$
4. Icterids are unusual in songbirds because they have considerable sexual dimorphism. $16.9$ $14.5$
@@
5. Couches are usually bought in a set together with cushions, which give them a bouncey and decorative comfortable touch. $-7.1$ $-6.8$
6. No, its 10 feet tall, and it’s a red big monster, demon like. $-7.3$ $-5.9$
7. It does this by following the grammatical basic rules of syntax. $-5.8$ $-6.0$
8. The daily average attendance from January through November 2010 was 22,133 people a day. $-5.8$ $-5.7$
@@
9. Being a spherical 3-manifold, it is the only homology 3-sphere, besides the 3-sphere itself, with a finite fundamental group. $0.9$ $16.8$
10. Since the 2015 Styria municipal structural reform, it is part of the municipality Birkfeld. $-0.5$ $15.3$
11. I won it at a fair last night. $0.2$ $15.4$
12. I say that the man who does not play whist lays up a sad old age for himself. $2.6$ $17.5$
@@
13. This is an ability which may have been present in their last common ancestor in the Archaean. $12.3$ $4.0$
14. Jacob, you’ve had a complex partial seizure, which can cause psychosis, including religious psychosis. $7.3$ $0.3$
15. He was also a right-arm medium fast bowler with three wickets in test matches to his credit. $6.2$ $-0.6$
16. He wheeled round sharply, and distinguished her lying with helpless outspread arms on the couch. $4.0$ $-2.8$
@@
A.2 Correlations to Cognitive Predictors
Correlations during training of AOP-$\Delta$ scores (Pythia-1.4b), with scores derived from various cognitive predictors.

### v2

@@ Results
  2. 2.
    Semantic cues: Non-immediate context can play a role in AOP under a different mechanism. Our results highlight the role of adjectives and nouns that occur quite early in the sentence. In the case of adjectives this is strongly driven by semantic similarity: top-scoring adjectives in the context have on average a cosine similarity of 0.15 with the first adjective in the phrase, and only 0.07 with the second adjective. [^4] Cosine similarity is computed using the input embeddings of the LM, summing subword embeddings where necessary.
6.3 Conclusions
The results in this section show that adjective order behaviour in LMs is context-sensitive. Building on the aggregate effects reported earlier (Figure 3C), the attribution analysis identifies two classes of contextual cues that shape model preferences: local collocational cues in the immediate vicinity of the adjective phrase, and earlier semantic cues that bias adjective ordering based on semantic similarity.
Importantly, these findings do not constitute an explanatory account of adjective order in human language. Rather, they illustrate how language models can serve as a tool for probing which aspects of sentence context are informative for adjective order preferences. In doing so, this analysis highlights contextual factors that are largely absent from existing theories of adjective order, and may serve as a starting point for future empirical studies on the role of context in human adjective order preferences.
@@ 7 Discussion
Generalization and Memorization in AOP
In this paper, we investigated the extent to which language models generalize adjective order preferences beyond direct memorization of training data statistics. Our results suggest that while models exhibit some degree of generalization, this generalization is limited and coexists with a strong reliance on distributional information. Here, we want to highlight an important question that is rarely brought up in linguistic generalization discussions in NLP: How general and abstract is human linguistic behaviour?
The linguistic literature suggests that human language production, including word order preferences, is guided by a mix of memorization and abstract principles (O’Donnell, 2015). Speakers tend to recycle more routinised and therefore more accessible combinations (MacDonald, 2013), guided by sequence frequencies in linguistic input that, in turn, form sequential associations that lead to production automatizing (Bybee, 2010; Diessel, 2019). This is compatible with a picture of language acquisition where knowledge of grammar is formed based on encountered exemplars of usage, which are stored in memory (Bybee, 2010) and serve as basis of generalization (Goldberg, 2006), but can also serve as patterns for routinised language production. So, human linguistic knowledge and language production is not totally abstract and is not based exclusively on generalization. This has been discussed across linguistic phenomena and has been suggested for word order (MacDonald, 2013; Levshina et al., 2023), although not for adjective order specifically.
When investigating generalization in LMs it is therefore important to make a distinction between the degree to which they create generalizations for a particular phenomenon, and how similar this generalization is to human behaviour. In principle, one could even find cases of LM overgeneralization. Importantly, in order to answer how human-like a model’s generalization behaviour is, we would need data on the trade-off in memorization and generalization in human linguistic behaviour: an empirical question for which detailed human data on adjective order preferences is currently lacking.
Corpus Interventions
Our analysis of AOP generalization is constrained by the fixed training corpus the models are trained on. More detailed insights into this question can be obtained by an intervention-based approach, in which we filter out specific constructions from the training data (Jumelet et al., 2021; Misra and Mahowald, 2024; Patil et al., 2024; Müller-Eberstein et al., 2025). An interesting experiment in this direction would be to remove all sentences containing double adjectives from the training corpus, and measure whether the remaining distribution of adjectives contains a sufficient signal to acquire adjective order constraints.
Cross-lingual Extensions
Our analysis is limited to adjective order preferences in English, which exhibits ordering patterns that differ from those found in other languages. Cross-linguistic research has documented both variation and regularities in adjective ordering, motivating general theories that go beyond any single language (Cinque, 2010; Culbertson et al., 2020). We leave a systematic cross-linguistic investigation of adjective order preferences in language models to future work.
@@
Despite the impressive performance of LMs in recent years, their impact on theoretical linguistics has remained minimal (Baroni, 2022). There are various reasons for this. On the one hand, the cognitive plausibility of current LMs remains questionable, both in terms of the scale at which they are trained (Huebner et al., 2021; Wilcox et al., 2025) and their inductive biases that depend strongly on model architecture and are not reflective of human language processing (Wilson and Frank, 2023; Oh et al., 2024; Padovani et al., 2025). On the other hand, their black box nature makes it highly challenging to extract linguistic patterns that can be tested against existing linguistic theories.
In this paper, we treat language models not as explanatory theories of adjective order, but as empirical tools for analysing how soft word order constraints are acquired by LMs. Although language models substantially outperform individual cognitive predictors proposed in prior work, predictive accuracy alone does not constitute an explanation of adjective order preferences in human language. Instead, we view their value as lying in the combination of behavioural analysis with interpretability methods, which makes it possible to probe what sources of information influence model decisions (Futrell and Mahowald, 2025). In particular, the attribution-based analysis in § 6 illustrates how such methods can be used to identify contextual cues that systematically modulate adjective order preferences—factors that are largely absent from existing accounts and that may serve as targets for future empirical investigation in human language.
Acknowledgements
We gratefully thank the anonymous reviewers for their feedback on an earlier version of the paper, which greatly helped in shaping the story. We thank Kanishka Misra, Kyle Mahowald and Jennifer Culbertson for the inspiration that our adjacent collaboration has provided for finalizing this paper. Jaap Jumelet is supported by NWO grant VI.Vidi.221C.009.
@@ Appendix A Complementary Results
A.1 CAP Examples
We provide a sample of CAP examples in Table A.1, categorised by the contextual and isolated AOP scores of Pythia-12b.
CAP Sentence AOP()($\bullet$) AOP(|c)($\bullet$|c)
black       CAP order always preferred
1. The seat comprises the following electoral wards $14.5$ $23.6$
2. The local wildlife includes leopards, deer, monkeys, several kinds of pheasants, the marbled polecat and the increasingly rare flying squirrel $11.1$ $24.9$
3. He was elected as leader of the Labour Party in the first round of voting using the single transferable vote $13.4$ $21.7$
4. Hoggett and mutton can taste more flavorful than lamb because they contain a higher concentration of species- characteristic fatty acids $12.8$ $22.1$
5. Jimmy seemed ready to take up that challenge, and entered upon an argument calculated to prove that he was a mild mannered individual $14.4$ $18.4$
@@ A.1 CAP Examples
6. Mm, well, the extraction of foreign rectal bodies 11.3-11.3 11.3-11.3
7. did they sing it’s an old grand flag 7.7-7.7 14.1-14.1
8. His great jaw protruded frightfully, with the upper thin lip 6.6-6.6 6.4-6.4
9. And Stuart sat down in, because they got these wooden old chairs 6.9-6.9 6.0-6.0
10. It does this by following the grammatical basic rules 5.9-5.9 6.1-6.1
@@
11. Be it thy course to busy giddy minds $1.9$ $26.2$
12. I have not passed through fire and death to bandy crooked words $5.3$ $29.2$
13. "I wish to preach, not the doctrine of ignoble ease, but the doctrine of the strenuous life; the life of toil and effort; of labor and strife; to preach that highest form of success which comes, not to the man who desires mere easy peace $2.3$ $21.8$
14. The "Yonsei" are the subject of on- going academic research 0.2-0.2 $18.1$
15. The city flag of Portland, Oregon, consists of a green field on which is placed a white 4-pointed star from which radiate blue stripes, each bordered by L- shaped yellow lines 1.0-1.0 $16.9$
@@
16. One expected that it earned a alone decisive battle $9.1$ 1.1-1.1
17. Meager is known for its last major eruption $12.8$ $4.8$
18. As a matter of fact, no body in the universe revolves about the exact geometrical center $7.6$ 0.4-0.4
19. E major is a musical major scale $4.8$ 2.8-2.8
20. That worthy was seated before a table spread with papers, and as Tom entered or rather was pushed into his presence he compressed his beetling black brows $5.1$ 2.5-2.5
@@
A.2 Overlap and Correlations to Cognitive Predictors
Overlap between various cognitive predictors and AOP-$\Delta$ scores of Pythia-1.4b::
[Uncaptioned image][Uncaptioned image]
Correlations during training of AOP-$\Delta$ scores (Pythia-1.4b), with scores derived from various cognitive predictors: [Uncaptioned image]
@@ A.5 Context Improvement Ratios
A.6 Context Length Impact
Relation between context length and contextual AOP-$\Delta$ scores, averaged over randomly sampled contexts. The context length effect shown in Figure 7 disappears when averaging over random contexts, showing that contextual gain is not driven by context length itself.